lucille is 4 times as old as nicole, six years from now, she will be twicw as old as nicole,how old are they no?
Answer by TimothyLamb(4379)

x = lucy age now
y = nick age now
---
x = 4y
x + 6 = 2(y + 6)
---
put the system of linear equations into standard form
---
x - 4y = 0
x + 6 = 2y + 12
---
x - 4y = 0
x - 2y = 6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = lucy age now = 12
y = nick age now = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php